The tocloft package^{∗}

Peter Wilson Catholic University of America^{†}

             Now at peter.r.wilson@boeing.com

1999/03/07

Abstract

The tocloft package provides means of controlling the typographic design of the Table of Contents, List of Figures and List of Tables. Use of this package requires the stdclsdv package.

This is a beta release. It has been tested with the tocbibind and float packages, and maybe others. Later versions may not necessarily be upwards compatible.

Contents

1 Introduction 1 1.1 L^{A}TEX’s methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 The tocloft package 4 2.1 Changing the titles . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Typesetting the entries . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.3 Experimental utilities . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.4 Usage with other packages . . . . . . . . . . . . . . . . . . . . . . . 9

3 The package code 10

List of Figures

1 Layout of a ToC (LoF, LoT) entry . . . . . . . . . . . 2

1 Introduction

In the standard classes the typographic design of the Table of Contents (ToC), the List of Figures (LoF) and List of Tables (LoT) is fixed or, more precisely, it is buried within the class definitions. The tocloft package provides handles for an author to change the design to meet the needs of the particular document.

∗This file has version number v0.2b, last revised 1999/03/07.

†This work was originally performed as a Guest Researcher at the National Institute of Standards and Technology.

[figure: \linewidth; \@pnumwidth; \@tocrmarg; indent; numwidth; 487; title end; continue…; … title; 3.5; Heading…; … title; .; .; .; .; .; .; .; .; .; .; .; .; \@dotsep]

Figure 1: Layout of a ToC (LoF, LoT) entry

Elements of the package were developed as part of a class and package bundle for typesetting ISO standards [Wil96]. This manual is typeset according to the conventions of the L^{A}TEX docstrip utility which enables the automatic extraction of the L^{A}TEX macro source files [GMS94].

Section 2 describes the usage of the package. Commented source code for the package is in Section 3. The package requires the stdclsdv package.

This is the first beta release of the package; if possible, later versions will be upwards compatible. It has been tested in combination with at least the tocbibind package and the algorithm package (which, in turn, calls the float package). Please send me any comments as to how you think that the package can be improved, or of any interesting examples of how you have used it.^{1}

1.1 L^{A}TEX’s methods

This is a general description of how L^{A}TEX does the processing for a Table of Contents. As the processing for List of Figures and List of Tables is similar I will, without loss of generality, just discuss the ToC.

\addcontentsline L^{A}TEX generates a .toc file if the document contains a \tableofcontents command. The sectioning commands^{2} put entries into the .toc file by calling the L^{A}TEX \addcontentsline{⟨file⟩}{⟨kind⟩}{⟨title⟩} command, where ⟨file⟩ is the file extension (e.g., toc), ⟨kind⟩ is the kind of entry (e.g., section or subsection), and ⟨title⟩ is the (numberered) title text. In the cases where there is a number, the ⟨title⟩ argument is given in the form {\numberline{number} title-text}.

\contentsline The \addcontentsline command writes an entry to the given file in the form \contentsline{⟨kind⟩}{⟨title⟩}{⟨page⟩} where ⟨page⟩ is the page number. For each ⟨kind⟩, L^{A}TEX provides a command \l@kind{⟨title⟩}{⟨page⟩} which performs the actual typesetting of the \contentsline entry.

\@pnumwidth The general layout of a typeset entry is illustrated in Figure 1. There are three

\@tocrmarg
\@dotsep

internal L^{A}TEX commands that are used in the typesetting. The page number is typeset flushright in a box of width \@pnumwidth, and the box is at the righthand margin. If the page number is too long to fit into the box it will stick out into the righthand margin. The title text is indented from the righthand margin by an amount given by \@tocrmarg. Note that \@tocrmarg should be greater than \@pnumwidth. Some entries are typeset with a dotted leader between the end of the title title text and the righthand margin indentation. The distance, in math units^{3} between the dots in the leader is given by the value of \@dotsep. In the standard classes the same values are used for the ToC, LoF and the LoT.

1Thanks to Rowland (rebecca@astrid.u-net.com), John Foster (john@isjf.demon.co.uk), Kasper (kbg@dkik.dk), Lee Nave (nave@math.washington.edu), and Andrew Thurber (athurber@emba.uvm.edu) for their suggestions.

2For figures and tables it is the \caption command that populates the .lof and .lot files.

The standard values for these internal commands are:

   • \@pnumwidth = 1.55em
   • \@tocrmarg = 2.55em
   • \@dotsep = 4.5

The values can be changed by using \renewcommand, in spite of the fact that the first two appear to be lengths.

Dotted leaders are not available for Part and Chapter ToC entries (nor for Section entries in the article class and its derivatives).

\numberline Each \l@kind macro is responsible for setting the general indent from the lefthand margin, and the numwidth. The \numberline{⟨number⟩} macro is responsible for typesetting the number flushleft in a box of width numwidth. If the number is too long for the box then it will protrude into the title text. The title text is indented by (indent + numwidth) from the lefthand margin. That is, the title text is typeset in a block of width

(\linewidth - indent - numwidth - \@tocrmarg).
   You can use the \addcontentsline command to add \contentsline com-

mands to a file.

\addtocontents L^{A}TEX also provides the \addtocontents{⟨file⟩}{⟨text⟩} command that will insert ⟨text⟩ into ⟨file⟩. You can use this for adding extra text and/or macros into the file, for processing when the file is typeset by \tableofcontents (or whatever other command is used for ⟨file⟩ processing, such as \listoftables for a .lot file).

As \addcontentsline and addtocontents write their arguments to a file, any fragile commands used in their arguments must be \protected.

You can make certain adjustments to the ToC etc., layout without using any package. Some examples are:

• If your page numbers stick out into the righthand margin

       \renewcommand{\@pnumwidth}{3em} \renewcommand{\@tocrmarg}{4em}

but using lengths appropriate to your document.

• The dots in the leaders can be eliminated by increasing \@dotsep to a large value:

       \renewcommand{\@dotsep}{10000}

3There are 18mu to 1em.

• To have dotted leaders in your ToC and LoF but not in your LoT:

 ...
 \tableofcontents
 \makeatletter \renewcommand{\@dotsep}{10000} \makeatother
 \listoftables
 \makeatletter \renewcommand{\@dotsep}{4.5} \makeatother
 \listoffigures
 ...

For this document I used this method to double the dot spacing for the LoF with respect to that for the ToC. As you can see, it is much better that all dot leaders have the same spacing.

• To add a horizontal line across the whole width of the ToC below an entry for a Part:

 \part{Part title}
 \addtocontents{toc}{\protect\mbox{}\protect\hrulefill\par}

Note that as both \addtocontents and \addcontentsline write their ar-

guments to a file, it means that any fragile commands in their arguments must be protected by preceeding each fragile command with \protect. The result of the example above would be the following two lines in the .toc file (assuming that it is the second Part and is on page 34):

 \contentsline {part}{II\hspace {1em}Part title}{34}
 \mbox {}\hrulefill \par

If the \protects were not used, then the second line would instead be:

 \unhbox \voidb@x \hbox {}\unhbox \voidb@x \leaders \hrule \hfill \kern \z@ \par

Remember, if you are modifying any command that includes an @ sign then this must be done in either a .sty file or if in the document itself it must be surrounded by \makeatletter and \makeatother. For example, if you want to modify \@dotsep in the preamble to your document you have to do it like this:

\makeatletter
\renewcommand{\@dotsep}{9.0}
\makeatother

2 The tocloft package

The tocloft package provides means of specifying the typography of the Table of Contents (ToC), the List of Figures (LoF) and the List of Tables (LoT).

\tableofcontents The ToC, LoF, and LoT are printed at the point in the document where these

\listoffigures
\listoftables

commands are called, as per normal L^{A}TEX. However, there is one difference between the standard L^{A}TEX behaviour and the behaviour with the tocloft package. In the standard L^{A}TEX classes that have \chapter headings, the ToC, LoF and LoT each appear on a new page. With the tocloft package they do not automatically start new pages; if you want them to be on new pages you have to specifically issue an appropriate command beforehand. For example:

  ...
 \clearpage
 \tableofcontents
 \clearpage
 \listoftables
 ...

2.1 Changing the titles

Commands are provided for controlling the appearance of the titles. Following L^{A}TEX custom, the title texts are the values of the \contentsname,

\listfigurename and \listtablename commands.

Similar sets of commands are provided for ToC, LoF and LoT title typsetting control. For convenience (certainly mine, and hopefully yours) in the following descriptions I will use ‘X’ to stand for ‘toc’ or ‘lof’ or ‘lot’. For example, \cftmarkX

stands for \cftmarktoc or \cftmarklof or \cftmarklot.

\cftmarkX These macros set the appearance of the running heads on the ToC, LoF, and

LoT pages. You probably don’t need to change these.

\cftbeforeXtitleskip These lengths control the vertical spacing before and after the titles. You can \cftafterXtitleskip change them from their default values by using \setlength.

\cftXtitlefont The code used for typesetting the ToC title looks like

\cftafterXtitle
 {\cfttoctitlefont \contentsname}{\cftaftertoctitle}\par

By default, \cftXtitlefont is defined as a font specification (e.g., \Large\bfseries), and \cftafterXtitle is empty. These commands can be changed (via \renewcommand) to change the typesetting. As examples:

   • \renewcommand{\cftXtitlefont}{\hfill\Large\itshape} will result in

a Large italic title typeset flushright.

   • \renewcommand{\cftXtitlefont}{\hfill\Large\bfseries} together with
     \renewcommand{\cftafterXtitle}{\hfill} will give a centered Large

bold title.

• Doing

      \renewcommand{\cftafterXtitle}{%
        \\[\baselineskip]\mbox{}\hfill{\normalfont Page}}

will put the word ‘Page’ flushright on the line following the title. (If you do this, then you may need to decrease \cftafterXtitleskip).

2.2 Typesetting the entries

Commands are also provided to enable finer control over the typesetting of the different kinds of entries. The parameters defining the default layout of the entries are illustrated as part of the layouts package or in [GMS94, page 34], and are repeated in Figure 1.

\cftdot In the default ToC typsetting only the more minor entries have dotted leader lines between the sectioning title and the page number. The tocloft package provides for general leaders for all entries. The ‘dot’ in a leader is given by the value of \cftdot. Its default definition is \newcommand{\cftdot}{.} which gives the default dotted leader. By changing \cftdot you can use symbols other than a period in the leader. For example

 \renewcommand{\cftdot}{\ensuremath{\ast}}

will result in a dotted leader using asterisks as the symbol.

\cftdotsep Each kind of entry can control the seperation between the dots in its leader \cftnodots (see below). For consistency though, all dotted leaders should use the same spac-

ing. The macro \cftdotsep specifies the default spacing. Its value is a number. However, if the seperation is too large then no dots will be actually typeset. The macro \cftnodots is a seperation value that is ‘too large’.

\cftsetpnumwidth The page numbers are typeset in a fixed width box. The command

\cftsetrmarg \cftsetpnumwidth{⟨length⟩} can be used to change the width of the box (L^{A}TEX’s internal \@pnumwidth). The title texts will end before reaching the righthand margin. \cftsetrmarg{⟨length⟩} can be used to set this distance (L^{A}TEX’s internal \@tocrmarg). Note that the length used in \cftsetrmarg should be greater than the length set in \cftsetpnumwidth. These values should remain constant in any given document.

In the following I will use X to stand for the following:

• part for \part titles

   • chap for \chapter titles
   • sec for \section titles
   • subsec for \subsection titles
   • subsubsec for \subsubsection titles
   • para for \paragraph titles
   • subpara for \subparagraph titles

• fig for figure \caption titles

• tab for table \caption titles

\cftbeforeXskip This controls the vertical space before an entry. It can be changed by using

\setlength.

\cftXindent This controls the indentation of an entry from the left margin (indent in Figure 1). It can be changed using \setlength. NOTE: This command is not provided for part as their entries will have no indentation (they are the highest level document divisions).

\cftXnumwidth This controls the space allowed for typesetting title numbers (numwidth in Figure 1). It can be changed using \setlength. Second and subsequent lines of a multiline title will be indented by this amount.

The remaining commands are related to the specifics of typesetting an entry. This is a simplified version of the code for typesetting numbered entry and unnumbered entries.

 {\cftXfont SNUM\cftXaftersnum \cftXaftersnumb TITLE}%
         {\cftXleader}{\cftXpagefont PAGE}\cftXafterpnum\par
 {\cftXfont TITLE}{\cftXleader}{\cftXpagefont PAGE}\cftXafterpnum\par

where SNUM is the section number, TITLE is the title text and PAGE is the page number. Note that for a Part entry the unnumbered style is used, whether or not the Part has a number.

\cftXfont This controls the appearance of the title (and its preceeding number, if any).

It may be changed using \renewcommand.

\cftXaftersnum The section number is typeset within a box and the \cftXaftersnum macro is \cftXaftersnumb called after the number is typeset but before the box is completed. After the box

is typeset the \cftXaftersnumb macro is called before typesetting the title text. Both of these can be changed by \renewcommand. By default they are defined to do nothing.

\cftXleader \cftXleader defines the leader between the title and the page number; it \cftXdotsep can be changed by \renewcommand. The spacing between any dots in the leader

is controlled by \cftXdotsep (\@dotsep in Figure 1). It can be changed by \renewcommand and its value must be either a number (e.g., 6.6 or \cftdotsep) or \cftnodots (to disable the dots). The spacing is in terms of math units where there are 18mu to 1em.

\cftXpagefont This defines the font to be used for typesetting the page number. It can be

changed by \renewcommand.

\cftXafterpnum This macro is called after the page number has been typeset. Its default is to

do nothing. It can be changed by \renewcommand.

Various effects can be achieved by changing the definitions of \cftXfont,

\cftXaftersnum, \cftXaftersnumb, \cftXleader and \cftXafterpnum, either

singly or in combination. For the sake of some examples, assume that we have the following initial definitions

 \newcommand{\cftXfont}{\bfseries}
 \newcommand{\cftXaftersnum}{}
 \newcommand{\cftXaftersnumb}{}
 \newcommand{\cftXleader}{\bfseries\@cftdotfill{\cftXdotsep}}
 \newcommand{\cftXdotsep}{\cftdotsep}
 \newcommand{\cftXpagefont}{\bfseries}
 \newcommand{\cftXafterpnum}{}

(Note that the same font should be used for the title, leader and page number to provide a coherent appearance).

• To eliminate the dots in the leader:

      \renewcommand{\cftXdotsep}{\cftnodots}

• To put something (e.g., a name) before the title (number):

      \renewcommand{\cftXfont}{\bfseries SOMETHING }

• To add a colon after the section number:

      \renewcommand{\cftXaftersnum}{:}

• To put something before the title number, add a colon after the the title number, and start the title text on the following line:

      \renewcommand{\cftXfont}{\bfseries SOMETHING }
      \renewcommand{\cftXaftersnum}{:}
      \renewcommand{\cftXaftersnumb}{\\}

• To set the title ragged left (but this only looks good for single line titles):

      \renewcommand{\cftXfont}{\hfill\bfseries}
      \renewcommand{\cftXleader}{}

• To set the page number immediately after the title instead of at the righthand margin:

      \renewcommand{\cftXleader}{}
      \renewcommand{\cftXafterpnum}{\cftparfillskip}

By default the \parfillskip value is locally set to fill up the last line of a paragraph. Just changing \cftXleader puts horrible interword spaces into the last line of the title. The \cftparfillskip command^{4} is part of the tocloft package and is provided just so that the above effect can be achieved.

At this point, I leave it up to your ingenuity as to other effects that you can achieve. However, if you come up with further examples, let me know for possible inclusion in a later version of this document.

2.3 Experimental utilities

The macros described in this section are even more experimental than those described previously.

\cftchapterprecis Some old style novels, and even some modern text books,^{5} include a short synopsis of the contents of the chapter either immediately after the chapter heading or in the Toc, or in both places.

The command \cftchapterprecis{⟨text⟩} prints its argument both at the point in the document where it is called, and also adds it to the .toc file. For example:

4Is there a better name for this?

5For example, Robert Sedgewick, Algorithms, Addison-Wesley, 1983.

 ...
 \chapter{} % first chapter
 \cftchapterprecis{Our hero is introduced; family tree; early days.}
 ...

\cftchapterprecishere The \cftchapterprecis command calls these two commands to print the text \cftchapterprecistoc in the document (the \...here{⟨text⟩} command) and to put it into the ToC (the

\...toc{⟨text⟩} command). These can be used individually if required.

Sometimes it may be desirable to make a change to the global parameters for an individual entry. For example, a figure might be placed on the end paper of a book (the inside of the front or back cover), and this needs to be placed in a LoF with the page number set as, say ‘inside front cover’. If ‘inside front cover’ is typeset as an ordinary page number it will stick out into the margin. Therefore, the parameters for this particular entry need to be changed.

\cftlocalchange The command \cftlocalchange{⟨file⟩}{⟨pnumwidth⟩}{⟨tocrmarg⟩} will write an entry into ⟨file⟩ to reset the global parameters. The command should be called again after any special entry to reset the parameters back to their usual values. Any fragile commands used in the arguments must be protected.

\cftaddtitleline The command \cftaddtitleline{⟨file⟩}{⟨kind⟩}{⟨title⟩}{⟨page⟩} will write a \contentsline entry into ⟨file⟩ for a ⟨kind⟩ entry with title ⟨title⟩ and page number ⟨page⟩. That is, an entry is made of the form:

\contentsline{kind}{title}{page}

Any fragile commands used in the arguments must be protected.

\cftaddnumtitleline  The command \cftaddnumtitleline{⟨file⟩}{⟨kind⟩}{⟨num⟩}{⟨title⟩}{⟨page⟩}

is similar except that it also includes ⟨num⟩ as the argument to the \numberline. That is, an entry is made of the form:

\contentsline{kind}{\numberline{num} title}{page}

Any fragile commands used in the arguments must be protected.

As an example of the use of these commands, noting that the default L^{A}TEX values for \@pnumwidth and \@tocrmarg are 1.55em and 2.55em respectively, one might do the following for a figure on the frontispiece page.

 ...
 % this is the frontispiece page with no number
 % draw or import the picture (with no \caption)
 \cftlocalchange{lof}{4em}{5em} % make pnumwidth big enough for
                                % frontispiece and change margin to suit
 \cftaddtitleline{lof}{figure}{The title}{frontispiece}
 \cftlocalchange{lof}{1.55em}{2.55em} % return to normal settings
 ...

Recall that a \caption command will put an entry in the .lof file, which is not wanted here. If a caption is required, then you can either craft one youself or, assuming that your general captions are not too exotic, use the \legend command from the ccaption package. If the illustration is numbered, use the

\cftaddnumtitleline command instead of \cftaddtitleline.

2.4 Usage with other packages

The tocloft and tocbibind packages can be used together in the same document. The tocbibind package provides easy means of adding document elements like the bibliography or the index to the Table of Contents. However there are two known potential problems:

• The 1998/11/15 version of tocbibind may give surprising results if the

     \toctocname, \toclotname or \toclofname commands have been used.

You should consider getting the current version of tocbibind.

• If the argument to the \tocextrahead command is other than one of the normal sectioning divisions (i.e., part through to sub-paragraph) such as \tocextrahead{clause}, then this will almost certainly cause a problem (as the tocloft package will not know how to define the corresponding \l@clause command). In such a case you will have to supply the appropriate macros youself.

\@cftasnum Some packages, like the float package by Anselm Lingnau, enable the creation of \@cftasnumb other kinds of List of …. The tocloft package is only minimally able to change the

formatting of these, principally because the packages are independent of each other and, in the case of the float package, new kinds of float environments and their associated lists can be created on the fly at any point in a document. Some aspects of the typesetting are controlled by \@cftasnum and \@cftasnumb commands. These are equivalent to the \cftXaftersnum and \cftXaftersnumb commands described earlier. By default they are defined to do nothing, but may be renewed to do something.

In general, use with other packages that redefine any of the macros that tocloft also modifies is likely to be problematic.

3 The package code

Announce the name and version of the package, which requires L^{A}TEX 2ε and the stdclsdv package.

1 ⟨∗usc⟩

 2 \NeedsTeXFormat{LaTeX2e}
 3 \ProvidesPackage{tocloft}[1999/03/07 v0.2b parameterised ToC, etc., typesetting]
 4 \RequirePackage{stdclsdv}

In order to try and avoid name clashes with other packages, each internal name will include the character string @cft.

The stdclsdv package is used to determine what sectioning commands may be used in the document. We assume that if the class is known and it has \chapter commands, then it is either a book of report class; if it has \section commands then it is an article class or one of its derivatives.

Issue a warning if there are no recognised sectional divisions and then skip the rest of the package code.

 5 \ifSCDnodivs
 6 \PackageWarning{tocloft}%
 7  {I don't recognize any sectional divisions so I'll do nothing}
 8 \renewcommand{\SCDquit}{\endinput}
 9 \fi

Perhaps quit now.

10 \SCDquit

Issue a warning if the class is unknown. Use chapter style if \chapter is defined, else use the \section style if it is defined, otherwise quit.

11 \ifSCDknownclass\else
12 \PackageWarning{tocloft}%
13  {I don't recognize the class but I'll do my best}
14 \ifSCDnodivs
15  \PackageWarning{tocloft}%
16  {I don't recognize any sectional divisions so I'll do nothing}
17  \renewcommand{\SCDquit}{\endinput}
18 \else
19  \ifSCDchapter
20  \PackageWarning{tocloft}{The document class has chapter divisions}
21  \else
22  \ifSCDsection
23  \PackageWarning{tocloft}{The document class has section divisions}
24  \else
25  \PackageWarning{tocloft}%
26  {The class has neither chapters nor sections, so I'll do nothing}
27  \renewcommand{\SCDquit}{\endinput}
28  \fi
29  \fi
30 \fi
31 \fi
32 \SCDquit

\if@cfttocbibind A flag that is set TRUE iff the tocbibind package has been loaded. The 1998/11/15 version of tocbibind does not necessarily work well with tocloft.

33 \newif\if@cfttocbibind
34 \AtBeginDocument{%
35 \@ifpackageloaded{tocbibind}{\@cfttocbibindtrue}{\@cfttocbibindfalse}
36 \if@cfttocbibind
37  \@ifpackagelater{tocbibind}{1998/11/16}{}{%
38  \PackageWarning{tocloft}{%
39 You are using a version of the tocbibind package\MessageBreak
40 that is not compatible with tocloft.\MessageBreak
41 The results may be surprising.\MessageBreak
42 Consider installing the current version of tocbibind.}}
43 \fi

44 }

\cftmarktoc \cftmarklof \cftmarklot These three macros set the style for running heads. They are initialised to give the default appearance.

45 \newcommand{\cftmarktoc}{%
46 \@mkboth{\MakeUppercase\contentsname}{\MakeUppercase\contentsname}}
47 \newcommand{\cftmarklof}{%
48 \@mkboth{\MakeUppercase\listfigurename}{\MakeUppercase\listfigurename}}
49 \newcommand{\cftmarklot}{%
50 \@mkboth{\MakeUppercase\listtablename}{\MakeUppercase\listtablename}}

\@cfttocstart \@cfttocfinish Two macros to perform the actions at the beginning and end of the \tableofcontents command (and friends). \@cfttocstart deals with chaptered documents, ensuring that the ToC is typeset in a single column (see classes.dtx for the original code).

51 \newcommand{\@cfttocstart}{%
52 \ifSCDchapter
53  \if@twocolumn
54  \@restonecoltrue\onecolumn
55  \else
56  \@restonecolfalse
57  \fi
58 \fi}

\@cfttocfinish resets, if required, twocolumn typesetting.

59 \newcommand{\@cfttocfinish}{%
60 \ifSCDchapter
61  \if@restonecol\twocolumn\fi
62 \fi}

\@cftdobibtoc If the tocbibind package has been used and it has redefined \tableofcontents we need to cater for that. The contents of the definition are defined in tocbibind.

63 \newcommand{\@cftdobibtoc}{%
64 \if@dotoctoc
65  \if@bibchapter
66  \addcontentsline{toc}{chapter}{\contentsname}
67  \else
68  \addcontentsline{toc}{\@tocextra}{\contentsname}
69  \fi
70 \fi}

\tableofcontents This is a parameterised version of the default \tableofcontents command. Each class has its own definition, but we have to cater for all classes in one definition, hence some of the checks. The definition is modified after all packages have been loaded.

71 \AtBeginDocument{%
72 \renewcommand{\tableofcontents}{%
73 \@cfttocstart

Ensure that any previous paragraph has been finished. Within a group set the local paragraphing style and typeset the title.

74 \par
75 \begingroup
76  \parindent\z@ \parskip\z@
77  \@cftmaketoctitle

If tocbibind has been used, then add the ToC name to the ToC.

78  \if@cfttocbibind
79  \@cftdobibtoc
80  \fi

Finally, read the .toc file and finish up.

81  \@starttoc{toc}%
82 \endgroup
83 \@cfttocfinish}

84 }

\@cftmaketoctitle This command typesets the title for the ToC.

85 \newcommand{\@cftmaketoctitle}{%

86 \vspace*{\cftbeforetoctitleskip}
87 \interlinepenalty\@M
88 {\cfttoctitlefont\contentsname}{\cftaftertoctitle}
89 \cftmarktoc
90 \par\nobreak
91 \vskip \cftaftertoctitleskip
92 \@afterheading}

\cftbeforetoctitleskip \cftaftertoctitleskip These two lengths control the vertical spacing before and after the ToC title.

93 \newlength{\cftbeforetoctitleskip}
94 \newlength{\cftaftertoctitleskip}

Their values depend on whether the document has chapters or not. In chaptered documents the default ToC title is typeset as a \chapter*, otherwise as a

\section*.
95 \ifSCDchapter
96 \setlength{\cftbeforetoctitleskip}{50pt}
97 \setlength{\cftaftertoctitleskip}{40pt}
98 \else
99 \setlength{\cftbeforetoctitleskip}{3.5ex \@plus 1ex \@minus .2ex}
100 \setlength{\cftaftertoctitleskip}{2.3ex \@plus.2ex}
101 \fi
\cfttoctitlefont \cftaftertoctitle The ToC title is typeset in the style given by \cfttoctitlefont. The macro

\cftaftertoctitle is called after typesetting the title. This is initialised to do nothing. Both these macros can be redefined to do other things (e.g., adding an \hfill to \cfttoctitlefont will make the title flushright).

102 \ifSCDchapter
103 \newcommand{\cfttoctitlefont}{\normalfont\Huge\bfseries}
104 \else
105 \newcommand{\cfttoctitlefont}{\normalfont\Large\bfseries}
106 \fi
107 \newcommand{\cftaftertoctitle}{}
\cftsetpnumwidth \cftsetrmarg Users commands for setting \@pnumwidth and \@tocrmarg.
108 \newcommand{\cftsetpnumwidth}[1]{\renewcommand{\@pnumwidth}{#1}}
109 \newcommand{\cftsetrmarg}[1]{\renewcommand{\@tocrmarg}{#1}}

\cftdot \@cftdotfill In the default ToC, a dotted line can be used to provide a leader between a title and the page number. The definition of this leader is buried in the \@dottedtocline command. The \@cftdotfill{⟨sep⟩} command provides a parameterised version of the leader code, where ⟨sep⟩ is the seperation between the dots in mu units. The symbol used for the ‘dots’ in the leader is given by the value of \cftdot.

110 \newcommand{\cftdot}{.}
111 \newcommand{\@cftdotfill}[1]{%
112 \leaders\hbox{$\m@th\mkern #1 mu\hbox{\cftdot}\mkern #1 mu$}\hfill}

\cftdotsep \cftnodots \cftdotsep holds the default dot seperation. If the kerns in \@cftdotfill are large enough, then no dots will be printed. \cftnodots should be ‘large enough’.

113 \newcommand{\cftdotsep}{4.5}
114 \newcommand{\cftnodots}{10000}

Now for the trickier bits regarding the typesetting of the ToC entries.

A .toc (also .lof and .lot) file consists of a list of \contentsline{⟨kind⟩}{⟨title⟩}{⟨page⟩} commands, where ⟨kind⟩ is the kind of heading (e.g., part or section or figure), ⟨title⟩ is the title text (including the number), and ⟨page⟩ is the page number. The entries are inserted into the file by calling the \addcontentsline{⟨file⟩}{⟨kind⟩}{⟨title⟩} command, where ⟨file⟩ is the file extension (e.g., toc, lot) and the other arguments are the same as for the \contentsline command. (Arbitrary stuff may also be put into the file via the \addtocontents{⟨file⟩}{⟨text⟩} command). The typesetting of the \contentsline entries is performed by commands of the form \l@kind. The sectioning and captioning commands call \addcontentsline to insert their titles into the .toc etc., files.

For the purposes at hand it is generally impossible to teat the typesetting of a title and its number seperately, as both are bundled into the ⟨title⟩ argument within \contentsline. They could be handled seperately if the \contentsline command was suitably modified. If this was done, then the \addtocontentsline command would also need to be changed which would then require the sectioning and captioning commands to be modified as well. This is certainly possible, but would cause problems if any other package also modified the sectioning or captioning commands, and there are several packages which do this.

Having said this, for all but Part entries, the sectional number is typeset via the \numberline command. We can take advantage of this fact.

I have taken the decision to not touch the \contentsline macro and instead to do what can be done with it as it exists. That is, I will modify the \l@kind commands. Essentially, my new definitions consist of inlined versions of the code

for \@dottedtocline.
\cftparfillskip The \l@kind commands modify (locally) the value of \parfillskip. \cftparfillskip

is a copy of the default TEXbook \parfillskip definition.

115 \newcommand{\cftparfillskip}{\parfillskip=0pt plus1fil}

\numberline The purpose of the \numberline{⟨secnum⟩} command is to typeset ⟨secnum⟩ left justified in a box of width \@tempdima. I redefine it to add two additional parameters, namely \@cftasnum and \@cftasnumb (see ltsect.dtx for the original definition).

116 \renewcommand{\numberline}[1]{%
117 \hb@xt@\@tempdima{#1\@cftasnum\hfil}\@cftasnumb}

\@cftasnum \@cftasnumb Originally these were not defined but were \let to appropriate commands in the \l@... commands, but they have to be defined in case something unexpected calls \numberline, for example through use of the float package.^{6}

118 \newcommand{\@cftasnum}{}
119 \newcommand{\@cftasnumb}{}

\l@part \if@cftdopart \l@part{⟨title⟩}{⟨page⟩} typesets the ToC entry for a part heading. It is a parameterised copy of the default \l@part (see classes.dtx for the original definition and the code below for \l@subsection for an explanation of most of this code). By default, Parts (and Chapters) do not have dotted leaders. This package provides for all entries to have dotted leaders.

6This bug was discovered by Andrew Thurber when using the tocloft and algorithm packages together.

120 \newif\if@cftdopart
121 \ifSCDpart
122 \renewcommand*{\l@part}[2]{%
123 \@cftdopartfalse
124 \ifnum \c@tocdepth >-2\relax
125  \ifSCDchapter
126  \@cftdoparttrue
127  \fi
128  \ifnum \c@tocdepth >\m@ne
129  \ifSCDchapter\else
130  \@cftdoparttrue
131  \fi
132  \fi
133 \fi
134 \if@cftdopart
135  \ifSCDchapter
136  \addpenalty{-\@highpenalty}%
137  \else
138  \addpenalty\@secpenalty
139  \fi
140  \addvspace{\cftbeforepartskip}%
141  \@tempdima \cftpartnumwidth\relax
142  \let\@cftasnum \cftpartaftersnum
143  \let\@cftasnumb \cftpartaftersnumb
144  \begingroup
145  \parindent \z@ \rightskip \@pnumwidth
146  \parfillskip -\@pnumwidth
147  {\leavevmode
148  {\cftpartfont #1}{\cftpartleader}
149  {\hb@xt@\@pnumwidth{\hss {\cftpartpagefont #2}}}\cftpartafterpnum}\par
150  \nobreak
151  \ifSCDchapter
152  \global\@nobreaktrue
153  \everypar{\global\@nobreakfalse\everypar{}}%
154  \else
155  \if@compatibility
156  \global\@nobreaktrue
157  \everypar{\global\@nobreakfalse\everypar{}}%
158  \fi
159  \fi
160  \endgroup
161 \fi}
162 \fi
\cftbeforepartskip \cftpartnumwidth \cftpartfont \cftpartaftersnum \cftpartaftersnumb \cftpartleader \cftpartdotsep \cftpartpagefont \cftpartafterpnum These are the user commands to control the typesetting of Part entries. They are

initialised to give the standard appearance.

163 \newlength{\cftbeforepartskip}
164 \setlength{\cftbeforepartskip}{2.25em \@plus\p@}
165 \newlength{\cftpartnumwidth}
166 \setlength{\cftpartnumwidth}{3em}
167 \newcommand{\cftpartfont}{\large\bfseries}
168 \newcommand{\cftpartaftersnum}{}
169 \newcommand{\cftpartaftersnumb}{}
170 \newcommand{\cftpartleader}{\large\bfseries\@cftdotfill{\cftpartdotsep}}

171 \newcommand{\cftpartdotsep}{\cftnodots}
172 \newcommand{\cftpartpagefont}{\large\bfseries}
173 \newcommand{\cftpartafterpnum}{}

\l@chapter \l@chapter{⟨title⟩}{⟨page⟩} typesets the ToC entry for a chapter heading. It is a parameterised copy of the default \l@chapter (see classes.dtx for the original definition). This only applies to chaptered documents.

174 \ifSCDchapter
175 \renewcommand*{\l@chapter}[2]{%
176 \ifnum \c@tocdepth >\m@ne
177  \addpenalty{-\@highpenalty}%
178  \vskip \cftbeforechapskip
179  {\leftskip \cftchapindent\relax
180  \rightskip \@tocrmarg
181  \parfillskip -\rightskip
182  \parindent \cftchapindent\relax\@afterindenttrue
183  \interlinepenalty\@M
184  \leavevmode
185  \@tempdima \cftchapnumwidth\relax
186  \let\@cftasnum \cftchapaftersnum
187  \let\@cftasnumb \cftchapaftersnumb
188  \advance\leftskip \@tempdima \null\nobreak\hskip -\leftskip
189  {\cftchapfont #1}\nobreak
190  {\cftchapleader}
191  \nobreak
192  \hb@xt@\@pnumwidth{\hfil\cftchappagefont #2}\cftchapafterpnum\par}%
193 \fi}
194 \fi
\cftbeforechapskip \cftchapindent \cftchapnumwidth \cftchapfont \cftchapaftersnum \cftchapaftersnumb \cftchapleader \cftchapdotsep \cftchappagefont \cftchapafterpnum These are the user commands to control the typesetting of Chapter entries. They

are initialised to give the standard appearance.

195 \newlength{\cftbeforechapskip}
196 \setlength{\cftbeforechapskip}{1.0em \@plus\p@}
197 \newlength{\cftchapindent}
198 \setlength{\cftchapindent}{0em}
199 \newlength{\cftchapnumwidth}
200 \setlength{\cftchapnumwidth}{1.5em}
201 \newcommand{\cftchapfont}{\bfseries}
202 \newcommand{\cftchapaftersnum}{}
203 \newcommand{\cftchapaftersnumb}{}
204 \newcommand{\cftchapleader}{\bfseries\@cftdotfill{\cftchapdotsep}}
205 \newcommand{\cftchapdotsep}{\cftnodots}
206 \newcommand{\cftchappagefont}{\bfseries}
207 \newcommand{\cftchapafterpnum}{}

\l@section \l@section{⟨title⟩}{⟨page⟩} typesets the ToC entry for a section heading. It is a parameterised copy of the default \l@section (see classes.dtx for the original definition).

208 \renewcommand*{\l@section}[2]{%
209 \ifnum \c@tocdepth >\z@
210  \ifSCDchapter
211  \vskip \cftbeforesecskip
212  \else

213  \addpenalty\@secpenalty
214  \addvspace{\cftbeforesecskip}
215  \fi
216  {\leftskip \cftsecindent\relax
217  \rightskip \@tocrmarg
218  \parfillskip -\rightskip
219  \parindent \cftsecindent\relax\@afterindenttrue
220  \interlinepenalty\@M
221  \leavevmode
222  \@tempdima \cftsecnumwidth\relax
223  \let\@cftasnum \cftsecaftersnum
224  \let\@cftasnumb \cftsecaftersnumb
225  \advance\leftskip \@tempdima \null\nobreak\hskip -\leftskip
226  {\cftsecfont #1}\nobreak
227  {\cftsecleader}
228  \nobreak
229  \hb@xt@\@pnumwidth{\hfil\cftsecpagefont #2}\cftsecafterpnum\par}%
230 \fi}
\cftbeforesecskip \cftsecindent \cftsecnumwidth \cftsecfont \cftsecaftersnum \cftsecaftersnumb \cftsecleader \cftsecdotsep \cftsecpagefont \cftsecafterpnum These are the user commands to control the typesetting of Section entries. They

are initialised to give the standard appearance.

231 \newlength{\cftbeforesecskip}
232 \ifSCDchapter
233 \setlength{\cftbeforesecskip}{\z@ \@plus.2\p@}
234 \else
235 \setlength{\cftbeforesecskip}{1.0em \@plus\p@}
236 \fi
237 \newlength{\cftsecindent}
238 \ifSCDchapter
239 \setlength{\cftsecindent}{1.5em}
240 \else
241 \setlength{\cftsecindent}{0em}
242 \fi
243 \newlength{\cftsecnumwidth}
244 \ifSCDchapter
245 \setlength{\cftsecnumwidth}{2.3em}
246 \else
247 \setlength{\cftsecnumwidth}{1.5em}
248 \fi
249 \ifSCDchapter
250 \newcommand{\cftsecfont}{\normalfont}
251 \else
252 \newcommand{\cftsecfont}{\bfseries}
253 \fi
254 \newcommand{\cftsecaftersnum}{}
255 \newcommand{\cftsecaftersnumb}{}
256 \ifSCDchapter
257 \newcommand{\cftsecleader}{\normalfont\@cftdotfill{\cftsecdotsep}}
258 \else
259 \newcommand{\cftsecleader}{\bfseries\@cftdotfill{\cftsecdotsep}}
260 \fi
261 \ifSCDchapter
262 \newcommand{\cftsecdotsep}{\cftdotsep}
263 \else

264 \newcommand{\cftsecdotsep}{\cftnodots}
265 \fi
266 \ifSCDchapter
267 \newcommand{\cftsecpagefont}{\normalfont}
268 \else
269 \newcommand{\cftsecpagefont}{\bfseries}
270 \fi
271 \newcommand{\cftsecafterpnum}{}

\l@subsection \l@subsection{⟨title⟩}{⟨page⟩} typesets the ToC entry for a subsection heading. It is a parameterised copy of the default \l@subsection (see classes.dtx for the original definition).

272 \renewcommand*{\l@subsection}[2]{%

Only typeset the entry if it falls within the tocdepth.

273 \ifnum \c@tocdepth >\@ne

Add some vertical space.

274  \vskip \cftbeforesubsecskip

Start a group to keep paragraphing changes local. Set the \leftskip to the entry’s indentation.

275  {\leftskip \cftsubsecindent\relax

Set the \rightskip to \@tocrmarg to leave room for the page number.

276  \rightskip \@tocrmarg

Ensure that the last line of the entry will be filled. Setting \parfillskip to a negative number prevents any overfull box messages.

277  \parfillskip -\rightskip

Set the paragraph indent to the entry’s indentation.

278  \parindent \cftsubsecindent\relax\@afterindenttrue

Try and prevent breaks between lines in a multiple line entry.

279  \interlinepenalty\@M

Make sure that we have left vertical mode.

280  \leavevmode

Our version of \numberline expects that the width of the number box is in \@tempdima, and that the two macros \@cftasnum and \@cftasnumb are defined. We set all these to the values for this entry.

281  \@tempdima \cftsubsecnumwidth\relax
282  \let\@cftasnum \cftsubsecaftersnum
283  \let\@cftasnumb \cftsubsecaftersnumb

Arrange that the (section number and) first line of the title is set at the current indent, and any further lines are further indented.

284  \advance\leftskip \@tempdima \null\nobreak\hskip -\leftskip

Print the (number and) title, prohibiting any breaking.

285  {\cftsubsecfont #1}\nobreak

Print the leader between the title and the page number, again asking for no breaks.

286  {\cftsubsecleader}
287  \nobreak

Finally, set the page number flushright in a box of width \@pnumwidth, not forgetting to finish the paragraph and close the group.

288  \hb@xt@\@pnumwidth{\hfil\cftsubsecpagefont #2}\cftsubsecafterpnum\par}%
289 \fi}
\cftbeforesubsecskip \cftsubsecindent \cftsubsecnumwidth \cftsubsecfont \cftsubsecaftersnum \cftsubsecaftersnumb \cftsubsecleader \cftsubsecdotsep \cftsubsecpagefont \cftsubsecafterpnum These are the user commands to control the typesetting of Sub-section entries.

They are initialised to give the standard appearance.

290 \newlength{\cftbeforesubsecskip}
291 \setlength{\cftbeforesubsecskip}{\z@ \@plus.2\p@}
292 \newlength{\cftsubsecindent}
293 \ifSCDchapter
294 \setlength{\cftsubsecindent}{3.8em}
295 \else
296 \setlength{\cftsubsecindent}{1.5em}
297 \fi
298 \newlength{\cftsubsecnumwidth}
299 \ifSCDchapter
300 \setlength{\cftsubsecnumwidth}{3.2em}
301 \else
302 \setlength{\cftsubsecnumwidth}{2.3em}
303 \fi
304 \newcommand{\cftsubsecfont}{\normalfont}
305 \newcommand{\cftsubsecaftersnum}{}
306 \newcommand{\cftsubsecaftersnumb}{}
307 \newcommand{\cftsubsecleader}{\normalfont\@cftdotfill{\cftsubsecdotsep}}
308 \newcommand{\cftsubsecdotsep}{\cftdotsep}
309 \newcommand{\cftsubsecpagefont}{\normalfont}
310 \newcommand{\cftsubsecafterpnum}{}
\l@subsubsection  \l@subsubsection{⟨title⟩}{⟨page⟩} typesets the ToC entry for a subsubsection

heading. It is a parameterised copy of the default \l@subsubsection (see classes.dtx for the original definition).

311 \renewcommand*{\l@subsubsection}[2]{%
312 \ifnum \c@tocdepth >\tw@
313  \vskip \cftbeforesubsubsecskip
314  {\leftskip \cftsubsubsecindent\relax
315  \rightskip \@tocrmarg
316  \parfillskip -\rightskip
317  \parindent \cftsubsubsecindent\relax\@afterindenttrue
318  \interlinepenalty\@M
319  \leavevmode
320  \@tempdima \cftsubsubsecnumwidth\relax
321  \let\@cftasnum \cftsubsubsecaftersnum
322  \let\@cftasnumb \cftsubsubsecaftersnumb
323  \advance\leftskip \@tempdima \null\nobreak\hskip -\leftskip
324  {\cftsubsubsecfont #1}\nobreak
325  {\cftsubsubsecleader}
326  \nobreak
327  \hb@xt@\@pnumwidth{\hfil\cftsubsubsecpagefont #2}\cftsubsubsecafterpnum\par}%
328 \fi}
\cftbeforesubsubsecskip \cftsubsubsecindent \cftsubsubsecnumwidth \cftsubsubsecfont \cftsubsubsecaftersnum \cftsubsubsecaftersnumb \cftsubsubsecleader \cftsubsubsecdotsep \cftsubsubsecpagefont \cftsubsubsecafterpnum These are the user commands to control the typesetting of Sub-sub-section entries.

They are initialised to give the standard appearance.

329 \newlength{\cftbeforesubsubsecskip}
330 \setlength{\cftbeforesubsubsecskip}{\z@ \@plus.2\p@}
331 \newlength{\cftsubsubsecindent}
332 \ifSCDchapter
333 \setlength{\cftsubsubsecindent}{7.0em}
334 \else
335 \setlength{\cftsubsubsecindent}{3.8em}
336 \fi
337 \newlength{\cftsubsubsecnumwidth}
338 \ifSCDchapter
339 \setlength{\cftsubsubsecnumwidth}{4.1em}
340 \else
341 \setlength{\cftsubsubsecnumwidth}{3.2em}
342 \fi
343 \newcommand{\cftsubsubsecfont}{\normalfont}
344 \newcommand{\cftsubsubsecaftersnum}{}
345 \newcommand{\cftsubsubsecaftersnumb}{}
346 \newcommand{\cftsubsubsecleader}{\normalfont\@cftdotfill{\cftsubsubsecdotsep}}
347 \newcommand{\cftsubsubsecdotsep}{\cftdotsep}
348 \newcommand{\cftsubsubsecpagefont}{\normalfont}
349 \newcommand{\cftsubsubsecafterpnum}{}

\l@paragraph \l@paragraph{⟨title⟩}{⟨page⟩} typesets the ToC entry for a paragraph heading. It is a parameterised copy of the default \l@paragraph (see classes.dtx for the original definition).

350 \renewcommand*{\l@paragraph}[2]{%
351 \ifnum \c@tocdepth >3\relax
352  \vskip \cftbeforeparaskip
353  {\leftskip \cftparaindent\relax
354  \rightskip \@tocrmarg
355  \parfillskip -\rightskip
356  \parindent \cftparaindent\relax\@afterindenttrue
357  \interlinepenalty\@M
358  \leavevmode
359  \@tempdima \cftparanumwidth\relax
360  \let\@cftasnum \cftparaaftersnum
361  \let\@cftasnumb \cftparaaftersnumb
362  \advance\leftskip \@tempdima \null\nobreak\hskip -\leftskip
363  {\cftparafont #1}\nobreak
364  {\cftparaleader}
365  \nobreak
366  \hb@xt@\@pnumwidth{\hfil\cftparapagefont #2}\cftparaafterpnum\par}%
367 \fi}
\cftbeforeparaskip \cftparaindent \cftparanumwidth \cftparafont \cftparaaftersnum \cftparaaftersnumb \cftparaleader \cftparadotsep \cftparapagefont \cftparaafterpnum These are the user commands to control the typesetting of Paragraph entries.

They are initialised to give the standard appearance.

368 \newlength{\cftbeforeparaskip}
369 \setlength{\cftbeforeparaskip}{\z@ \@plus.2\p@}
370 \newlength{\cftparaindent}
371 \ifSCDchapter
372 \setlength{\cftparaindent}{10em}
373 \else
374 \setlength{\cftparaindent}{7.0em}
375 \fi

376 \newlength{\cftparanumwidth}
377 \ifSCDchapter
378 \setlength{\cftparanumwidth}{5em}
379 \else
380 \setlength{\cftparanumwidth}{4.1em}
381 \fi
382 \newcommand{\cftparafont}{\normalfont}
383 \newcommand{\cftparaaftersnum}{}
384 \newcommand{\cftparaaftersnumb}{}
385 \newcommand{\cftparaleader}{\normalfont\@cftdotfill{\cftparadotsep}}
386 \newcommand{\cftparadotsep}{\cftdotsep}
387 \newcommand{\cftparapagefont}{\normalfont}
388 \newcommand{\cftparaafterpnum}{}
\l@subparagraph  \l@subparagraph{⟨title⟩}{⟨page⟩} typesets the ToC entry for a subparagraph

heading. It is a parameterised copy of the default \l@subparagraph (see classes.dtx for the original definition).

389 \renewcommand*{\l@subparagraph}[2]{%
390 \ifnum \c@tocdepth >4\relax
391  \vskip \cftbeforesubparaskip
392  {\leftskip \cftsubparaindent\relax
393  \rightskip \@tocrmarg
394  \parfillskip -\rightskip
395  \parindent \cftsubparaindent\relax\@afterindenttrue
396  \interlinepenalty\@M
397  \leavevmode
398  \@tempdima \cftsubparanumwidth\relax
399  \let\@cftasnum \cftsubparaaftersnum
400  \let\@cftasnumb \cftsubparaaftersnumb
401  \advance\leftskip \@tempdima \null\nobreak\hskip -\leftskip
402  {\cftsubparafont #1}\nobreak
403  {\cftsubparaleader}
404  \nobreak
405  \hb@xt@\@pnumwidth{\hfil\cftsubparapagefont #2}\cftsubparaafterpnum\par}%
406 \fi}
\cftbeforesubparaskip \cftsubparaindent \cftsubparanumwidth \cftsubparafont \cftsubparaaftersnum \cftsubparaaftersnumb \cftsubparaleader \cftsubparadotsep \cftsubparapagefont \cftsubparaafterpnum These are the user commands to control the typesetting of Sub-paragraph entries.

They are initialised to give the standard appearance.

407 \newlength{\cftbeforesubparaskip}
408 \setlength{\cftbeforesubparaskip}{\z@ \@plus.2\p@}
409 \newlength{\cftsubparaindent}
410 \ifSCDchapter
411 \setlength{\cftsubparaindent}{12em}
412 \else
413 \setlength{\cftsubparaindent}{10em}
414 \fi
415 \newlength{\cftsubparanumwidth}
416 \ifSCDchapter
417 \setlength{\cftsubparanumwidth}{6em}
418 \else
419 \setlength{\cftsubparanumwidth}{5em}
420 \fi
421 \newcommand{\cftsubparafont}{\normalfont}
422 \newcommand{\cftsubparaaftersnum}{}

423 \newcommand{\cftsubparaaftersnumb}{}
424 \newcommand{\cftsubparaleader}{\normalfont\@cftdotfill{\cftsubparadotsep}}
425 \newcommand{\cftsubparadotsep}{\cftdotsep}
426 \newcommand{\cftsubparapagefont}{\normalfont}
427 \newcommand{\cftsubparaafterpnum}{}

\@cftdobiblof If the tocbibind package has been used and it has redefined \listoffigures we need to cater for that. The contents of the definition are defined in tocbibind.

428 \newcommand{\@cftdobiblof}{%
429 \if@dotoclof
430  \if@bibchapter
431  \addcontentsline{toc}{chapter}{\listfigurename}
432  \else
433  \addcontentsline{toc}{\@tocextra}{\listfigurename}
434  \fi
435 \fi}

\listoffigures This is a parameterised version of the default \listoffigures command. The changes are postponed until after all packages have been loaded. Each class has its own definition, but we have to cater for all classes in one definition, hence some of the checks. First, perform the default checks for multicolumns.

436 \AtBeginDocument{
437 \renewcommand{\listoffigures}{%
438 \@cfttocstart

Ensure that any previous paragraph has been finished. Within a group set the local paragraphing style. Typeset the title and then do the contents of the .lof file.

439 \par
440 \begingroup
441  \parindent\z@ \parskip\z@
442  \@cftmakeloftitle
443  \if@cfttocbibind
444  \@cftdobiblof
445  \fi
446  \@starttoc{lof}%
447 \endgroup

Finally, restore any multicolumn typesetting.

448 \@cfttocfinish}

449 }

\@cftmakeloftitle This command typesets the title for the LoF.

450 \newcommand{\@cftmakeloftitle}{%
451 \vspace*{\cftbeforeloftitleskip}
452 \interlinepenalty\@M
453 {\cftloftitlefont\listfigurename}{\cftafterloftitle}
454 \cftmarklof
455 \par\nobreak
456 \vskip \cftafterloftitleskip
457 \@afterheading}

\cftbeforeloftitleskip \cftafterloftitleskip These two lengths control the vertical spacing before and after the LoF title.

458 \newlength{\cftbeforeloftitleskip}
459 \newlength{\cftafterloftitleskip}

Their values depend on whether the document has chapters or not. In chaptered documents the default LoF title is typeset as a \chapter*, otherwise as a

 \section*.
460 \ifSCDchapter
461 \setlength{\cftbeforeloftitleskip}{50pt}
462 \setlength{\cftafterloftitleskip}{40pt}
463 \else
464 \setlength{\cftbeforeloftitleskip}{3.5ex \@plus 1ex \@minus .2ex}
465 \setlength{\cftafterloftitleskip}{2.3ex \@plus.2ex}
466 \fi
\cftloftitlefont \cftafterloftitle The LoF title is typeset in the style given by \cftloftitlefont. The macro

\cftafterloftitle is called after typesetting the title. This is initialised to do nothing. Both these macros can be redefined to do other things (e.g., adding an \hfill to \cftloftitlefont will make the title flushright).

467 \ifSCDchapter
468 \newcommand{\cftloftitlefont}{\normalfont\Huge\bfseries}
469 \else
470 \newcommand{\cftloftitlefont}{\normalfont\Large\bfseries}
471 \fi
472 \newcommand{\cftafterloftitle}{}

\l@figure \l@figure{⟨title⟩}{⟨page⟩} typesets the LoF entry for a figure caption heading. It is a parameterised copy of the default \l@figure (see classes.dtx for the original definition).

473 \renewcommand*{\l@figure}[2]{%
474  \vskip \cftbeforefigskip
475  {\leftskip \cftfigindent\relax
476  \rightskip \@tocrmarg
477  \parfillskip -\rightskip
478  \parindent \cftfigindent\relax\@afterindenttrue
479  \interlinepenalty\@M
480  \leavevmode
481  \@tempdima \cftfignumwidth\relax
482  \let\@cftasnum \cftfigaftersnum
483  \let\@cftasnumb \cftfigaftersnumb
484  \advance\leftskip \@tempdima \null\nobreak\hskip -\leftskip
485  {\cftfigfont #1}\nobreak
486  {\cftfigleader}
487  \nobreak
488  \hb@xt@\@pnumwidth{\hfil\cftfigpagefont #2}\cftfigafterpnum\par}%

489 }

\cftbeforefigskip \cftfigindent \cftfignumwidth \cftfigfont \cftfigaftersnum \cftfigaftersnumb \cftfigleader \cftfigdotsep \cftfigpagefont \cftfigafterpnum These are the user commands to control the typesetting of Figure caption entries.

They are initialised to give the standard appearance.

490 \newlength{\cftbeforefigskip}
491 \setlength{\cftbeforefigskip}{\z@ \@plus.2\p@}
492 \newlength{\cftfigindent}
493 \setlength{\cftfigindent}{1.5em}
494 \newlength{\cftfignumwidth}
495 \setlength{\cftfignumwidth}{2.3em}
496 \newcommand{\cftfigfont}{\normalfont}
497 \newcommand{\cftfigaftersnum}{}

498 \newcommand{\cftfigaftersnumb}{}
499 \newcommand{\cftfigleader}{\normalfont\@cftdotfill{\cftfigdotsep}}
500 \newcommand{\cftfigdotsep}{\cftdotsep}
501 \newcommand{\cftfigpagefont}{\normalfont}
502 \newcommand{\cftfigafterpnum}{}

\@cftdobiblot If the tocbibind package has been used and it has redefined \listoftables we need to cater for that. The contents of the definition are defined in tocbibind.

503 \newcommand{\@cftdobiblot}{%
504 \if@dotoclot
505  \if@bibchapter
506  \addcontentsline{toc}{chapter}{\listtablename}
507  \else
508  \addcontentsline{toc}{\@tocextra}{\listtablename}
509  \fi
510 \fi}

\listoftables This is a parameterised version of the default \listoftables command. The changes are postponed until after all packages have been loaded. Each class has its own definition, but we have to cater for all classes in one definition, hence some of the checks. First, perform the default checks for multicolumns.

511 \AtBeginDocument{
512 \renewcommand{\listoftables}{%
513 \@cfttocstart

Ensure that any previous paragraph has been finished. Within a group set the local paragraphing style. Typeset the title and then do the contents of the .lot file.

514 \par
515 \begingroup
516  \parindent\z@ \parskip\z@
517  \@cftmakelottitle
518  \if@cfttocbibind
519  \@cftdobiblot
520  \fi
521  \@starttoc{lot}%
522 \endgroup

Finally, restore any multicolumn typesetting.

523 \@cfttocfinish}

524 }

\@cftmakelottitle This command typesets the title for the LoT.

525 \newcommand{\@cftmakelottitle}{%
526 \vspace*{\cftbeforelottitleskip}
527 \interlinepenalty\@M
528 {\cftlottitlefont\listtablename}{\cftafterlottitle}
529 \cftmarklot
530 \par\nobreak
531 \vskip \cftafterlottitleskip
532 \@afterheading}

\cftbeforelottitleskip \cftafterlottitleskip These two lengths control the vertical spacing before and after the LoT title.

533 \newlength{\cftbeforelottitleskip}
534 \newlength{\cftafterlottitleskip}

Their values depend on whether the document has chapters or not. In chaptered documents the default LoT title is typeset as a \chapter*, otherwise as a

 \section*.
535 \ifSCDchapter
536 \setlength{\cftbeforelottitleskip}{50pt}
537 \setlength{\cftafterlottitleskip}{40pt}
538 \else
539 \setlength{\cftbeforelottitleskip}{3.5ex \@plus 1ex \@minus .2ex}
540 \setlength{\cftafterlottitleskip}{2.3ex \@plus.2ex}
541 \fi
\cftlottitlefont \cftafterlottitle The LoT title is typeset in the style given by \cftlottitlefont. The macro

\cftafterlottitle is called after typesetting the title. This is initialised to do nothing. Both these macros can be redefined to do other things (e.g., adding an \hfill to \cftlottitlefont will make the title flushright).

542 \ifSCDchapter
543 \newcommand{\cftlottitlefont}{\normalfont\Huge\bfseries}
544 \else
545 \newcommand{\cftlottitlefont}{\normalfont\Large\bfseries}
546 \fi
547 \newcommand{\cftafterlottitle}{}

\l@table \l@table{⟨title⟩}{⟨page⟩} typesets the LoT entry for a table caption heading. It is a parameterised copy of the default \l@table (see classes.dtx for the original definition).

548 \renewcommand*{\l@table}[2]{%
549  \vskip \cftbeforetabskip
550  {\leftskip \cfttabindent\relax
551  \rightskip \@tocrmarg
552  \parfillskip -\rightskip
553  \parindent \cfttabindent\relax\@afterindenttrue
554  \interlinepenalty\@M
555  \leavevmode
556  \@tempdima \cfttabnumwidth\relax
557  \let\@cftasnum \cfttabaftersnum
558  \let\@cftasnumb \cfttabaftersnumb
559  \advance\leftskip \@tempdima \null\nobreak\hskip -\leftskip
560  {\cfttabfont #1}\nobreak
561  {\cfttableader}
562  \nobreak
563  \hb@xt@\@pnumwidth{\hfil\cfttabpagefont #2}\cfttabafterpnum\par}%

564 }

\cftbeforetabskip \cfttabindent \cfttabnumwidth \cfttabfont \cfttabaftersnum \cfttabaftersnumb \cfttableader \cfttabdotsep \cfttabpagefont \cfttabafterpnum These are the user commands to control the typesetting of Table caption entries.

They are initialised to give the standard appearance.

565 \newlength{\cftbeforetabskip}
566 \setlength{\cftbeforetabskip}{\z@ \@plus.2\p@}
567 \newlength{\cfttabindent}
568 \setlength{\cfttabindent}{1.5em}
569 \newlength{\cfttabnumwidth}
570 \setlength{\cfttabnumwidth}{2.3em}
571 \newcommand{\cfttabfont}{\normalfont}
572 \newcommand{\cfttabaftersnum}{}

573 \newcommand{\cfttabaftersnumb}{}
574 \newcommand{\cfttableader}{\normalfont\@cftdotfill{\cfttabdotsep}}
575 \newcommand{\cfttabdotsep}{\cftdotsep}
576 \newcommand{\cfttabpagefont}{\normalfont}
577 \newcommand{\cfttabafterpnum}{}

\cftchapterprecis This is experimental. \cftchapterprecis{⟨text⟩} typesets ⟨text⟩ at the point where it is called, and also adds ⟨text⟩ to the .toc file. It is expects to be called immediately after a \chapter command.

578 \newcommand{\cftchapterprecis}[1]{%
579 \cftchapterprecishere{#1}
580 \cftchapterprecistoc{#1}}

\cftchapterprecishere \cftchapterprecishere{⟨text⟩} typesets ⟨text⟩. It expects to be called immediately after a \chapter command. First add some negative vertical space to move it closer to the chapter heading.

581 \newcommand{\cftchapterprecishere}[1]{%
582 \vspace*{-2\baselineskip}

Typeset its argument using italic font in a quote environment.

583 \begin{quote}\textit{#1}\end{quote}}
\cftchapterprecistoc  \cftchapterprecistoc{⟨text⟩} adds ⟨text⟩ to the .toc file. The ⟨text⟩ will be

typeset within the same margins as the the title text of a \chapter heading, using an italic font.

584 \newcommand{\cftchapterprecistoc}[1]{\addtocontents{toc}{%

Start a group to localize changes to the paragraphing. Set the left margin to the chapter indent plus the chapter number width.

585 {\leftskip \cftchapindent\relax
586 \advance\leftskip \cftchapnumwidth\relax

Set the right hand margin to \@tocrmarg.

587 \rightskip \@tocrmarg\relax

Typeset ⟨text⟩ using an italic font, then ensure that the paragraph is finished (to use the local skips). Finally close the group and we are done.

588 \textit{#1}\protect\par}}}

\cftlocalchange \cftmakelocalchange{⟨file⟩}{⟨pnumwidth⟩}{⟨tocrmarg⟩} makes an entry into ⟨file⟩ to change the \@pnumwidth and the \@tocrmarg values.

589 \newcommand{\cftlocalchange}[3]{%
590 \addtocontents{#1}{\protect\cftsetpnumwidth{#2} \protect\cftsetrmarg{#3}}}
\cftaddtitleline  \cftaddtitleline{⟨file⟩}{⟨kind⟩}{⟨title⟩}{⟨page⟩} adds a \contentsline entry

to ⟨file⟩ with the given information.

591 \newcommand{\cftaddtitleline}[4]{%
592 \addtocontents{#1}{\protect\contentsline{#2}{#3}{#4}}}
\cftaddnumtitleline  \cftaddtitleline{⟨file⟩}{⟨kind⟩}{⟨num⟩}{⟨title⟩}{⟨page⟩} adds a \contentsline

entry to ⟨file⟩ with the given information.

593 \newcommand{\cftaddnumtitleline}[5]{%
594 \addtocontents{#1}%
595  {\protect\contentsline{#2}{\protect\numberline{#3}{\protect\ignorespaces #4}}{#5}}}

The end of this package.

596 ⟨/usc⟩

References

[GMS94] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The LaTeX Companion. Addison-Wesley Publishing Company, 1994.

[Wil96] Peter R. Wilson. LaTeX for standards: The LaTeX package files user manual. NIST Report NISTIR, June 1996.

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

Symbols

\@afterheading . . . .

. . . . . 92, 457, 532

\@afterindenttrue .

. . . . . . . . 182, 219, 278, 317, 356, 395, 478, 553

\@cftasnum . 10, 117,

\underline{118}, 142, 186, 223, 282, 321, 360, 399, 482, 557

\@cftasnumb . 10, 117,

\underline{118}, 143, 187, 224, 283, 322, 361, 400, 483, 558

\@cftdobiblof . \underline{428}, 444
\@cftdobiblot . \underline{503}, 519
\@cftdobibtoc . . . \underline{63}, 79
\@cftdopartfalse . . 123
\@cftdoparttrue 126, 130
\@cftdotfill . . \underline{110},

170, 204, 257, 259, 307, 346, 385, 424, 499, 574

\@cftmakeloftitle .

. . . . . . . . 442, \underline{450}

\@cftmakelottitle .

. . . . . . . . 517, \underline{525}

\@cftmaketoctitle 77, \underline{85}
\@cfttocbibindfalse 35
\@cfttocbibindtrue . 35
\@cfttocfinish . . . .

. . \underline{51}, 83, 448, 523

\@cfttocstart . . . . .

. . \underline{51}, 73, 438, 513 \@dotsep . . . . . . . . . . 2

\@highpenalty . 136, 177
\@ifpackagelater . . 37
\@ifpackageloaded . 35
\@nobreakfalse 153, 157
\@nobreaktrue . 152, 156
\@pnumwidth . . . . . . .

… 2, 108, 145, 146, 149, 192, 229, 288, 327, 366, 405, 488, 563

\@restonecolfalse . 56
\@restonecoltrue . . 54
\@secpenalty . . 138, 213
\@starttoc 81, 446, 521
\@tocextra 68, 433, 508
\@tocrmarg . . . . . . .

2, 109, 180, 217, 276, 315, 354, 393, 476, 551, 587

A

\addcontentsline . .

…. 2, 66, 68, 431, 433, 506, 508

\addtocontents . . 3,

584, 590, 592, 594

\AtBeginDocument . .

. . 34, 71, 436, 511

B

\baselineskip . . . . . 582

C

                          \c@tocdepth . . . 124,

128, 176, 209, 273, 312, 351, 390

                          \cftaddnumtitleline

. . . . . . . . . 9, \underline{593}

                          \cftaddtitleline 9, \underline{591}
                          \cftafterloftitle .

. . . . . . . . 453, \underline{467}

                          \cftafterloftitleskip

. . . . . . . . 456, \underline{458}

                          \cftafterlottitle .

. . . . . . . . 528, \underline{542}

                          \cftafterlottitleskip

. . . . . . . . 531, \underline{533}

                          \cftaftertoctitle .

. . . . . . . . 88, \underline{102}

                          \cftaftertoctitleskip

. . . . . . . . . 91, \underline{93}

                          \cftafterXtitle . . . . 5
                          \cftafterXtitleskip . 5
                          \cftbeforechapskip .

. . . . . . . . 178, \underline{195}

                          \cftbeforefigskip .

. . . . . . . . 474, \underline{490}

                          \cftbeforeloftitleskip

. . . . . . . . 451, \underline{458}

                          \cftbeforelottitleskip

. . . . . . . . 526, \underline{533}

                          \cftbeforeparaskip .

. . . . . . . . 352, \underline{368}

                          \cftbeforepartskip .

. . . . . . . . 140, \underline{163}

\cftbeforesecskip .

…. 211, 214, \underline{231}

\cftbeforesubparaskip

. . . . . . . . 391, \underline{407}

\cftbeforesubsecskip

. . . . . . . . 274, \underline{290}

\cftbeforesubsubsecskip

. . . . . . . . 313, \underline{329}

\cftbeforetabskip .

. . . . . . . . 549, \underline{565}

\cftbeforetoctitleskip

. . . . . . . . . 86, \underline{93}

\cftbeforeXskip . . . . 6
\cftbeforeXtitleskip 5
\cftchapafterpnum .

. . . . . . . . 192, \underline{195}

\cftchapaftersnum .

. . . . . . . . 186, \underline{195}

\cftchapaftersnumb .

. . . . . . . . 187, \underline{195}

\cftchapdotsep . . . . \underline{195}
\cftchapfont . . 189, \underline{195}
\cftchapindent . . . .

. 179, 182, \underline{195}, 585

\cftchapleader 190, \underline{195}
\cftchapnumwidth . .

…. 185, \underline{195}, 586

\cftchappagefont . .

. . . . . . . . 192, \underline{195}

\cftchapterprecis 8, \underline{578}
\cftchapterprecishere

. . . . . . 8, 579, \underline{581}

\cftchapterprecistoc

. . . . . . 8, 580, \underline{584} \cftdot . . . . . . . . 5, \underline{110}

\cftdotsep . . 6, \underline{113},

262, 308, 347, 386, 425, 500, 575

\cftfigafterpnum . .

. . . . . . . . 488, \underline{490}

\cftfigaftersnum . .

. . . . . . . . 482, \underline{490}

\cftfigaftersnumb .

. . . . . . . . 483, \underline{490}

\cftfigdotsep . . . . . \underline{490}
\cftfigfont . . . 485, \underline{490}
\cftfigindent . . . . .

…. 475, 478, \underline{490}

\cftfigleader . 486, \underline{490}
\cftfignumwidth 481, \underline{490}
\cftfigpagefont 488, \underline{490}
\cftlocalchange . 9, \underline{589}
\cftloftitlefont . .

. . . . . . . . 453, \underline{467}

\cftlottitlefont . .

. . . . . . . . 528, \underline{542}

\cftmarklof . . . . \underline{45}, 454
\cftmarklot . . . . \underline{45}, 529
\cftmarktoc . . . . . \underline{45}, 89

\cftmarkX . . . . . . . . . 5

\cftnodots . . . . . 6,

\underline{113}, 171, 205, 264

\cftparaafterpnum .

. . . . . . . . 366, \underline{368}

\cftparaaftersnum .

. . . . . . . . 360, \underline{368}

\cftparaaftersnumb .

. . . . . . . . 361, \underline{368}

\cftparadotsep . . . . \underline{368}
\cftparafont . . 363, \underline{368}
\cftparaindent . . . .

…. 353, 356, \underline{368}

\cftparaleader 364, \underline{368}
\cftparanumwidth . .

. . . . . . . . 359, \underline{368}

\cftparapagefont . .

. . . . . . . . 366, \underline{368}

\cftparfillskip . . . \underline{115}
\cftpartafterpnum .

. . . . . . . . 149, \underline{163}

\cftpartaftersnum .

. . . . . . . . 142, \underline{163}

\cftpartaftersnumb .

. . . . . . . . 143, \underline{163}

\cftpartdotsep . . . . \underline{163}
\cftpartfont . . 148, \underline{163}
\cftpartleader 148, \underline{163}
\cftpartnumwidth . .

. . . . . . . . 141, \underline{163}

\cftpartpagefont . .

. . . . . . . . 149, \underline{163}

\cftsecafterpnum . .

. . . . . . . . 229, \underline{231}

\cftsecaftersnum . .

. . . . . . . . 223, \underline{231}

\cftsecaftersnumb .

. . . . . . . . 224, \underline{231}

\cftsecdotsep . . . . . \underline{231}
\cftsecfont . . . 226, \underline{231}
\cftsecindent . . . . .

…. 216, 219, \underline{231}

\cftsecleader . 227, \underline{231}
\cftsecnumwidth 222, \underline{231}
\cftsecpagefont 229, \underline{231}
\cftsetpnumwidth . .

. . . . . . 6, \underline{108}, 590

\cftsetrmarg 6, \underline{108}, 590

                          \cftsubparaafterpnum

. . . . . . . . 405, \underline{407}

                          \cftsubparaaftersnum

. . . . . . . . 399, \underline{407}

                          \cftsubparaaftersnumb

. . . . . . . . 400, \underline{407}

                          \cftsubparadotsep . \underline{407}
                          \cftsubparafont 402, \underline{407}
                          \cftsubparaindent .

…. 392, 395, \underline{407}

                          \cftsubparaleader .

. . . . . . . . 403, \underline{407}

                          \cftsubparanumwidth

. . . . . . . . 398, \underline{407}

                          \cftsubparapagefont

. . . . . . . . 405, \underline{407}

                          \cftsubsecafterpnum

. . . . . . . . 288, \underline{290}

                          \cftsubsecaftersnum

. . . . . . . . 282, \underline{290}

                          \cftsubsecaftersnumb

. . . . . . . . 283, \underline{290}

                          \cftsubsecdotsep . . \underline{290}
                          \cftsubsecfont 285, \underline{290}
                          \cftsubsecindent . .

…. 275, 278, \underline{290}

                          \cftsubsecleader . .

. . . . . . . . 286, \underline{290}

                          \cftsubsecnumwidth .

. . . . . . . . 281, \underline{290}

                          \cftsubsecpagefont .

. . . . . . . . 288, \underline{290}

                          \cftsubsubsecafterpnum

. . . . . . . . 327, \underline{329}

                          \cftsubsubsecaftersnum

. . . . . . . . 321, \underline{329}

                          \cftsubsubsecaftersnumb

. . . . . . . . 322, \underline{329}

                          \cftsubsubsecdotsep  \underline{329}
                          \cftsubsubsecfont .

. . . . . . . . 324, \underline{329}

                          \cftsubsubsecindent

…. 314, 317, \underline{329}

                          \cftsubsubsecleader

. . . . . . . . 325, \underline{329}

                          \cftsubsubsecnumwidth

. . . . . . . . 320, \underline{329}

                          \cftsubsubsecpagefont

. . . . . . . . 327, \underline{329}

                          \cfttabafterpnum . .

. . . . . . . . 563, \underline{565}

                          \cfttabaftersnum . .

. . . . . . . . 557, \underline{565}

\cfttabaftersnumb .

. . . . . . . . 558, \underline{565}

\cfttabdotsep . . . . . \underline{565}
\cfttabfont . . . 560, \underline{565}
\cfttabindent . . . . .

…. 550, 553, \underline{565}

\cfttableader . 561, \underline{565}
\cfttabnumwidth 556, \underline{565}
\cfttabpagefont 563, \underline{565}
\cfttoctitlefont 88, \underline{102}
\cftXafterpnum . . . . . 7
\cftXaftersnum . . . . . 7
\cftXaftersnumb . . . . 7
\cftXdotsep . . . . . . . . 7

\cftXfont . . . . . . . . . 7

\cftXindent . . . . . . . . 6
\cftXleader . . . . . . . . 7
\cftXnumwidth . . . . . . 6
\cftXpagefont . . . . . . 7
\cftXtitlefont . . . . . 5
\contentsline 2, 592, 595
\contentsname . . . . .

…. 46, 66, 68, 88

E

\endinput …. 8, 17, 27

H

\hb@xt@ 117, 149, 192, 229, 288, 327, 366, 405, 488, 563

\Huge . . . . . 103, 468, 543

I

\if@bibchapter . . . .

. . . . . 65, 430, 505

\if@cftdopart . . . . . \underline{120}
\if@cfttocbibind . .

. . \underline{33}, 78, 443, 518

\if@compatibility . 155
\if@dotoclof . . . . . . 429
\if@dotoclot . . . . . . 504
\if@dotoctoc . . . . . . 64
\if@restonecol . . . . 61

                          \if@twocolumn . . . . . 53
                          \ifSCDchapter . . 19,

52, 60, 95, 102, 125, 129, 135, 151, 174, 210, 232, 238, 244, 249, 256, 261, 266, 293, 299, 332, 338, 371, 377, 410, 416, 460, 467, 535, 542

                          \ifSCDknownclass . . 11
                          \ifSCDnodivs . . . . . 5, 14
                          \ifSCDpart . . . . . . . 121
                          \ifSCDsection . . . . . 22
                          \ignorespaces . . . . . 595
                          \interlinepenalty 87,

183, 220, 279, 318, 357, 396, 452, 479, 527, 554

L

                          \l@chapter . . . . . . . \underline{174}

\l@figure . . . . . . . . \underline{473}

                          \l@paragraph . . . . . . \underline{350}

\l@part . . . . . . . . . . \underline{120}

                          \l@section . . . . . . . \underline{208}
                          \l@subparagraph . . . \underline{389}
                          \l@subsection . . . . . \underline{272}
                          \l@subsubsection . . \underline{311}

\l@table . . . . . . . . . \underline{548} \Large …. 105, 470, 545 \large …. 167, 170, 172 \leaders . . . . . . . . . 112

                          \leftskip . 179, 188,

216, 225, 275, 284, 314, 323, 353, 362, 392, 401, 475, 484, 550, 559, 585, 586

                          \listfigurename . . .

. 48, 431, 433, 453

                          \listoffigures . . 4, \underline{436}

                          \listoftables . . . 4, \underline{511}

                                                   \listtablename . . . .

. 50, 506, 508, 528

M

                                                   \MakeUppercase 46, 48, 50
                                                   \MessageBreak . . . 39–41

\mkern . . . . . . . . . . . 112

N \newif . . . . . . . . 33, 120

                                                   \normalfont 103, 105,

250, 257, 267, 304, 307, 309, 343, 346, 348, 382, 385, 387, 421, 424, 426, 468, 470, 496, 499, 501, 543, 545, 571, 574, 576

                                                   \numberline . . . . . 3, \underline{116}

O

                                                   \onecolumn . . . . . . . 54

P

                                                   \PackageWarning 6, 12,

15, 20, 23, 25, 38

                                                   \parindent . 76, 145,

182, 219, 278, 317, 356, 395, 441, 478, 516, 553

\parskip . . 76, 441, 516

                                                   \ProvidesPackage . . . 3

R

                                                   \RequirePackage . . . . 4

S \SCDquit 8, 10, 17, 27, 32

T

                                                   \tableofcontents . 4, \underline{71}

\textit . . . . . . 583, 588

                                                   \twocolumn . . . . . . . 61